how do you get the equation set up for the problem:
Kathy is six years younger than bill. 12 years ago, bill was twice as old as Kathy was then. how old are they now?

Answer by ankor@dixie-net.com(22740)

how do you get the equation set up for the problem:
Let k = Kathy's present age
Let b = Bill's present age
:
Write an equation for each statement simplify each equation:
:
"Kathy is six years younger than bill.
k = (b-6)
:
" 12 years ago, bill was twice as old as Kathy was then."
b - 12 = 2(k - 12)
b - 12 = 2k - 24
b - 12 = 2k - 24 + 12
b = 2k - 12
:
Now all you have to do is replace k with (b-6) and solve for b

how old are they now?